Sol Summary Report – March 3th

Crew 275 Sol Summary Report 03Mar2023

Sol: 19

Summary Title: Under the spotlight

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

Finally, a day when we wake up and don’t see the landscape covered with snow or mud! The flipside of this coin is that the temperature in the spacesuits got higher than usual and the EVA crew was very tired after their 4-hour EVA. We are no longer used to walk 10 000 steps a day, and even less in a spacesuit! An exhausting EVA, but a very fulfilling one, as all the objectives have been accomplished in perfect agreement with the planned timeline. The LOAC has been re-installed, while the batteries and the SD cards of all the other atmospheric instruments have been changed. All the checkpoints of the photogrammetry experiment have been identified by the subjects onsite at Candor Chasma, where the terrain was still slightly muddy.

The main event of the day was the presence of a journalist crew from BBC. They accompanied us during the first hour of our EVA. They then filmed the activities in the Upper Deck, the lunch, and finally they spent the whole afternoon interviewing us at our dedicated workplaces: the Science Dome, the GreenHab, etc. It was honestly a bit strange to see other people after three weeks just among us, but we are happy that they were very cooperative and respectful regarding the simulation. We are looking forward to seeing the output they can get out of their footage.

Look Ahead Plan:

Tomorrow (Sol 20), we would like to have our third EVA to Candor Chasma. The EVA crews will have to find specific checkpoints based on a 3D render. We would like to take the opportunity of this EVA to change the batteries of the atmospheric instruments, since they cannot last for more than two days, and we are not planning any EVA for Sunday.

Another media visit is planned for next Tuesday (Sol 23) with a crew from France Televisions. We will keep you updated as soon as we have more information regarding their planned arrival and departure.

Anomalies in work: None.

Weather: Cloudy, low -5°C/23°F, high 8°C/46°F.

Crew Physical Status: Optimal.

EVA: EVA # 15 was successful, both the main (photogrammetry) and secondary (atmospheric instruments) objectives have been fulfilled.

Reports to be filed: Sol Summary, EVA Report, EVA Request, Operations Report, Journalist Report, GreenHab Report, Astronomy Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site